# Flow of non-classical correlations in cluster states due to projective measurements

### v1

@@ I Introduction
Quantum information, unlike its classical counterpart, can lie delocalized across multiple physical systems. Quantum entanglement and more generic non-classical correlations that are quantified by measures like quantum discord are the underlying physical resources that allow quantum information to take on delocalized forms. The ability to store and manipulate information delocalized across a qubit register is a key difference between quantum and classical information processing. The significance of this difference becomes evident when the resources that enable quantum computers to perform certain tasks exponentially faster than the best-known classical approach are traced out. For instance, among the various quantum resources that have been identified as being useful for quantum information processing Chitambar and Gour (2019); Childs et al. (2018); Singh et al. (2015); Chitambar and Hsieh (2016); Veitch et al. (2014); Streltsov et al. (2017); Horodecki and Oppenheim (2012), it is known that for pure state quantum computing, multipartite entanglement that grows with the problem size is an essential requirement for exponential speedup Jozsa and Linden (2003). A comprehensive picture of the resources that facilitate a quantum advantage in information processing tasks beyond the pure state scenario is still not available but considerable progress has been made in this direction Bacon (2009); Gour (2024); Salazar et al. (2024); Berk et al. (2021); Lostaglio (2019); Bhattacharya et al. (2020).
@@
Following the dynamics of non-classical correlations while a quantum computation progresses is the focus of this Paper. Such dynamics, particularly that of quantum entanglement, has been considered in various contexts previously Życzkowski et al. (2001); Ptaszyński and Esposito (2024); Dahbi et al. (2022); Saghafi et al. (2019); Aolita et al. (2015). Here we look at it using cluster states and the measurement based quantum computation (MBQC) model. The MBQC model deviates from the conventional circuit model by initializing computations with cluster states as entangled resource states arranged on specific lattices or graphs, followed by the execution of local measurements Raussendorf and Briegel (2001); Raussendorf et al. (2003); Briegel et al. (2009). Local unitary operations on a single qubit cannot affect information delocalized across the multiple qubits that form an entangled cluster state. However, generic quantum operations that are not unitary applied locally on a single qubit can modify the non-classical correlations it shares with other qubits in the cluster. Measurements are a special class of such operations which are of interest to us in the following since it detaches the measured qubit from the rest of the cluster while at the same time modifying the delocalized information shared across the remaining qubits. We look at the dynamics of specific non-classical correlations that are shared between a few qubits of a cluster state as measurements are performed on the remaining qubits.
@@
The cluster state, which is the starting point of MBQC, is also a stabilizer state. From the Gottesman-Knill theorem it is known that stabilizer states, while being highly entangled, can also be simulated efficiently using classical means and so by themselves cannot yield quantum computational advantage Gottesman (1997); Aaronson and Gottesman (2004). Measurements on the individual qubits of the cluster not only remove the respective qubits from the cluster, they also can move the state of the remaining qubits away from being stabilizer states. These measurements modify the quantum information shared across all the qubits in a manner that allows for implementation of quantum algorithms that are exponentially faster than the best known classical ones on the cluster. Exploring how the delocalized information in cluster states respond to the measurements done on each of the qubits of the cluster is the subject matter of this Paper. We investigate this dynamics for simple, “ladder”-shaped cluster states and focus on the non-classical correlations that develop between two qubits in the last rung of the ladder as measurements are done on other qubits in the cluster.
@@
In Mahto et al. (2022) we investigated the relationship between non-classical correlations in a subsystem of a globally entangled state and genuine multipartite entanglement in the global state. We found that when the global state of a collection of qubits considered has a regular structure like a cluster state, the two are typically related. Curiously, precisely when the the global state is a stabilizer state, the subsystem discord vanishes even if the global state is highly entangled. We take this line of investigation forward here, but in a different direction. The ladder state is initially a stabilizer and the two qubits in the last rung share no quantum discord among them. We look at the non-classical correlations that develop between these two qubits as all the qubits in the cluster except the last three are measured. The measurements on the remaining qubits are chosen such that after each measurement, the state of the remaining qubits is pushed further away from being a stabilizer state. We choose to focus on the last three qubits of the ladder and study the quantum discord and entanglement between two of them because if the third one is also measured, the state of the remaining two will be pure and entanglement will be the only form of non-classical correlations present. With the third qubit also left unmeasured, there can be other forms of non-classical correlations between the last two.
@@
We can also view the last two qubits as an open quantum system coupled to a rather idealised environment that forms the rest of the ladder-like cluster state. From this perspective, we see how non-classical correlations can be built up in the open system due to measurements of a random nature happening on the environment that it is coupled to. As suggested in Mahto et al. (2022), the presence of non-classical correlations in a mixed state quantum computer can then be viewed as a consequence of computationally useful (non-stabilizer) entanglement present in the pure state of a larger system into which the mixed state of the computer is embedded in. The simple ladder-state model allows us to explore quantitatively how such correlations change in response to random measurements occurring on the environment.
@@
This paper is organized as follows. In Sec. II we review cluster states and their properties including non-classical correlations in subsystems. Flow of non-classical correlations in the ladder state model is discussed in the next section. Sec. IV contains further discussions and our conclusions.
@@
A cluster state Briegel and Raussendorf (2001) can be defined on a simple, un-directed graphs in any dimension. Qubits in the state $|+\rangle=(|0\rangle+|1\rangle)\sqrt{2}$ are placed on each of the vertices and each of the edges correspond to a two-qubit entangling gate $U_{kl}$ performed on the qubits placed on the vertices, $k$ and $l$ connected by the edge. The entangling gate used is typically the controlled-Z (CZ) gate. Cluster states are a special case of the more generic graph states Hartmann et al. (2007); Hein et al. (2006) wherein the initial states of the qubits on the vertices as well as the entangling operations done across the edges need not be uniform.
A cluster state can be constructed from a collection of qubits all in the state $|0\rangle$ through the application of the Hadamard and CZ gates only. Since these operations form a subset of the Clifford group Calderbank et al. (1998) of operations on a collection of qubits, the resultant cluster state is a stabilizer state Raussendorf et al. (2003). A stabilizer state $|\psi_{\rm stab}\rangle$ on $n$ qubits is the simultaneous eigenstate with $+1$ eigenvalue of $2^{n}$ operators, all of which are strings of Pauli operators (elements of the Clifford group). This allows for a succinct description of the state in terms of $n$ generators of the $2^{n}$ operators. Application of operators belonging to the Clifford group on such states results in states whose stabiliser generators can be computed directly from the generators of the initial state. In other words, Clifford group operations on stabilizer states produce other stabilizer states and any quantum information processing task restricted to Stabilizer states and Clifford group operations can be efficiently simulated using classical means, as proven in the celebrated Gottesman-Knill theorem Gottesman (1998); Aaronson and Gottesman (2004).
@@ II.1 Measurement based quantum computing
MBQC represents a distinctive approach to quantum computation that diverges from the traditional gate-based model. Gate-based quantum computing proceeds via application of quantum gates acting on a register of qubits, thereby evolving the state of the register towards a designated final state from which the output of the computation can be obtained Barenco et al. (1995); Nielsen and Chuang (2010). MBQC, however, takes a different path that rely on measurements for implementing the quantum computational paradigm. Instead of applying quantum gates directly, MBQC utilizes a cluster state as the substrate for its operations Nielsen (2006). Cluster states are an ideal starting point in that it provides a high degree of quantum entanglement to be used up as a resource as the computation proceeds. Given the graph corresponding to the desired cluster, the steps to be followed to prepare the cluster state is straightforward as well even if practical implementation can be challenging Kiesel et al. (2005); Zhang et al. (2006). Entanglement in a cluster state is remarkably resilient against local operations, making it more challenging to disrupt than other highly entangled states like Greenberger-Horne-Zeilinger (GHZ) states. In fact, a single local operation cannot annihilate all the entanglement in a multipartite cluster state, while the same is not true for a GHZ state Briegel and Raussendorf (2001); Gisin and Bechmann-Pasquinucci (1998).
MBQC is typically implemented on cluster states defined by simple and regular graphs like a two dimensional square lattice. Clusters defined on 2D triagonal, hexagonal and Kagome lattices are also known to be universal resource states for MBQC Van den Nest et al. (2006). In fact, highly entangled states that do not necessarily share the symmetry and robustness of entanglement of the clusters defined on these simple lattices are known to be not useful for MBQC Gross et al. (2009); Bremner et al. (2009). Staring from simple graphs like a 2D square lattice has the added advantage that given any quantum computation implemented through a gate-based quantum circuit, the same circuit can literally be “drawn” on a suitably large cluster state showing that both models are equivalent. Drawing such a circuit involves removal of several qubits from the cluster which is achieved by measuring all of them on to the eigenbasis of the Pauli-$Z$ operator (computational basis). Using simple and symmetric graphs for building the cluster is also helpful for writing down the stabilizer generators for the cluster. The $n$ generators are identified by associating a string of Pauli operators to each of the $n$ qubits in the lattice which consists of the Pauli-$X$ operator on the qubit along with $Z$ operators on all the qubits connected to it.
@@
Two measurement outcomes $s_{j}\in\{0,1\}$ are possible for each qubit and typically the measurement on the next qubit depends on the result of the measurement result on the previous one. Single qubit gates are implemented by measuring a row of connected qubits with suitable measurement angles $\varphi_{j}$. Up to five such measurements may be needed to implement an arbitrary single qubit rotation. A two qubit gate like the cnot gate can be implemented on the two-dimensional cluster through a similar sequence of measurements on two rows of qubits Raussendorf et al. (2003). In short, through the choice of suitable measurements a complete set of universal quantum gates can be realized within MBQC. Measurements on an $n$-qubit cluster state reproduces the action of a gate-based quantum computation on a register of $d$ qubits in the sense that after $n-d$ measurements, the state of the remaining qubits of the cluster will match the output of the gate-based circuit up to irrelevant overall phases. Measurement of these $d$ output qubits in a standard basis produces the measurement statistics which form the (classical) output of the computation.
@@ III Non-Classical Correlations in MBQC using Cluster States
The output qubits have to be driven to a specific state with its own distinctive quantum and classical correlations by the sequence of measurements in the MBQC model. Our interest is in how such correlations develop between the output qubits. For investigating this problem we consider a simplified cluster state defined on a square lattice with a ladder shape as shown in Fig. 1. The two qubits on the right end of the ladder are the output qubits. When the ladder shaped cluster is in the initial stabilizer state, non-classical correlations are not present in the state of the two output qubits Mahto et al. (2022). The question of interest is how such correlations develop as measurements are done on other qubits in the cluster leading to the localisation of the relevant information on these two qubits as the output of the computation. Do such correlations develop only when the immediate neighbours are measured or do they develop earlier indicating a flow of the quantum information delocalized across the cluster into these qubits at each stage? What is the nature of the non-classical correlations developed on the output qubits and how does it depend on the details of the measurements performed on the rest of the cluster. These are some of the questions that we investigate in the following.
@@
We quantify the non-classical correlations in $\rho_{34}^{\rm out}$ using quantum discord Ollivier and Zurek (2001); Henderson and Vedral (2001); Luo (2008); Modi et al. (2012); Bera et al. (2017). Quantum discord is essentially the difference between the total correlations and the classical correlations in a bipartite quantum state $\rho_{\rm AB}$ of two systems $A$ and $B$. It is an entropic measure and the total correlations are quantified in terms of the mutual information $I(A:B)=S(A)+S(B)+S(AB)$ between $A$ and $B$ with $S(X)=-{\rm tr}(\rho_{X}\log\rho_{X})$ denoting the vonNeumann entropy. Classical correlations are quantified in terms of the conditional entropy $S(A|\rho_{B}^{j})$ where $\rho_{B}^{j}$ is the post-measurement state of $B$ obtained when a measurement indexed by $j$ is performed on subsystem $B$. In terms of this, an alternate expression for the mutual information can be written down as $J(A:B)=S(A)-\sum_{j}S(A|\rho_{B}^{j})$, where the sum over $j$ ensures a complete set of measurements on $B$. If $A$ and $B$ are classical systems with the vonNeumann entropy replaced by the corresponding Shannon entropy, then by Bayes theorem, the two definitions of classical mutual information $I$ and $J$ are equal to each other. In the quantum case the two need not be equal. Defining ${\mathcal{J}}(A:\bar{B})=\max_{\{\Pi_{j}\}}J(A|B)$, where the maximum is over all possible measurements $\{\Pi_{j}\}$ on $B$, quantum discord is defined as the difference,
@@
We see that the two generators $g_{4}(2)$ and $g_{4}(4)$ remain Pauli strings, but the generator $g_{4}(3)$ is not a Pauli string anymore and the state of qubits $2$, $3$ and $4$ is no longer a stabilizer state. Since qubit $4$ is not directly connected to the measured qubit, the generator $g_{4}(4)$ associated with it remains invariant as well. At $\varphi_{1}=0,\pi/2,3\pi/2,$ and $2\pi$, the generator $g_{4}(3)$ also becomes a Pauli string and for these values of $\varphi_{1}$, the quantum discord $D_{3\bar{4}}$ is absent.
@@ III.1 Six and eight qubit ladders
Figure 4: The top panel shows $D_{5\bar{6}}$ as a function of $\varphi_{3}$ for 20,000 random choices of $(\varphi_{1},\varphi_{2},\varphi_{3})$ for the six qubit ladder state. The bottom panel shows $D_{\bar{5}6}$. Both $D_{5\bar{6}}$ and $D_{\bar{5}6}$ for each choice of $(\varphi_{1},\varphi_{2},\varphi_{3})$ are averaged over the eight possible measurement outcomes $(s_{j}=\pm 1)$.
We can track the flow of non-classical correlations in the six qubit ladder by computing $D_{\bar{5}6}$ when measurements are done one by one on qubits $1$, $2$ and $3$. When the measurements are on the first and second qubits only then both $D_{5\bar{6}}$ and $D_{\bar{5}6}$ are zero. This is because the measurement operators commute with the CZ gates applied between qubits (3,5), (4,6), and (5,6), thereby leaving the output state (5,6) unaffected. However, the effect of the measurements on $1$ and $2$ do eventually flow to qubits $5$ and $6$ when qubit $3$ is measured as can be seen from Fig. 4. As a function of the last measurement angle $\varphi_{3}$ shown in the bottom panel of Fig. 4, $D_{\bar{5}6}$ shows a discernible pattern with two distinct oscillatory patterns. As a function of the second-last measurement angle $\varphi_{3}$, the same pattern persists but is a bit less well defined. The top panel of Fig. 4 shows that even if there is a pattern in the relationship between $D_{\bar{5}6}$ and the first measurement angle $\varphi_{1}$, it is masked by the subsequent measurements and does not show up as clearly as in the other case. Note that the maximum values for $D_{5\bar{6}}$ for the six qubit case is same as the maximum for $D_{3\bar{4}}$ for four qubits. The stabilizer generators $g_{6}(4)$, $g_{6}(5)$ and $g_{6}(6)$ transform under the measurements specified by $(\varphi_{1},\varphi_{2},\varphi_{3})$ as
@@
Quantum discords $D_{7\bar{8}}$ and $D_{\bar{7}8}$ averaged over the $322^{5}=$ 32 possible measurement outcomes on the proceeding 5 qubits are plotted in Fig. LABEL:9qubitsa against the angle $\varphi_{5}$ of the measurement on qubit $5$ which is directly connected to qubit $7$ of the cluster. The discord is computed again for 20,000 randomly chosen values of the measurement angles $\varphi_{i}$, $i=1,\ldots,5$. We see that with the longer ladder, the non-classical correlations developed between the qubits in the last rung is stronger with both $D_{7\bar{8}}$ and $D_{\bar{7}8}$ having a maximum value of $1$. The asymmetry between the two discords also is not present indicating the presence of entanglement between the two. Entanglement as quantified by the concurrence was also computed for the 20,000 choices of measurement angles and the result is plotted in Fig. 7 showing that indeed the qubits in the last rung of the ladder does get entangled when the ladder is sufficiently long.

### v2

@@ I Introduction
Quantum information, unlike its classical counterpart, can lie delocalized across multiple physical systems. Quantum entanglement and more generic non-classical correlations that are quantified by measures like quantum discord are the underlying physical resources that allow quantum information to take on delocalized forms. The ability to store and manipulate information delocalized across a qubit register is a key difference between quantum and classical information processing. The significance of this difference becomes evident when the resources that enable quantum computers to perform certain tasks exponentially faster than the best-known classical approach are traced out. For instance, among the various quantum resources that have been identified as being useful for quantum information processing [1, 2, 3, 4, 5, 6, 7], it is known that for pure state quantum computing, multipartite entanglement that grows with the problem size is an essential requirement for exponential speedup [8]. A comprehensive picture of the resources that facilitate a quantum advantage in information processing tasks beyond the pure state scenario is still not available but considerable progress has been made in this direction [9, 10, 11, 12, 13, 14].
@@
Following the dynamics of non-classical correlations while a quantum computation progresses is the focus of this Paper. Such dynamics, particularly that of quantum entanglement, has been considered in various contexts previously [15, 16, 17, 18, 19]. Here we look at it using cluster states and the measurement based quantum computation (MBQC) model. The MBQC model deviates from the conventional circuit model by initializing computations with cluster states as entangled resource states arranged on specific lattices or graphs, followed by the execution of local measurements [20, 21, 22]. Local unitary operations on a single qubit cannot affect information delocalized across the multiple qubits that form an entangled cluster state. However, generic quantum operations that are not unitary applied locally on a single qubit can modify the non-classical correlations it shares with other qubits in the cluster. Measurements are a special class of such operations which are of interest to us in the following since it detaches the measured qubit from the rest of the cluster while at the same time modifying the delocalized information shared across the remaining qubits. We look at the dynamics of specific non-classical correlations that are shared between a few qubits of a cluster state as measurements are performed on the remaining qubits.
@@
The cluster state, which is the starting point of MBQC, is also a stabilizer state. From the Gottesman-Knill theorem it is known that stabilizer states, while being highly entangled, can also be simulated efficiently using classical means and so by themselves cannot yield quantum computational advantage [23, 24]. Measurements on the individual qubits of the cluster not only remove the respective qubits from the cluster, they also can move the state of the remaining qubits away from being stabilizer states. These measurements modify the quantum information shared across all the qubits in a manner that allows for implementation of quantum algorithms that are exponentially faster than the best known classical ones on the cluster. Exploring how the delocalized information in cluster states respond to the measurements done on each of the qubits of the cluster is the subject matter of this Paper. We investigate this dynamics for simple, “ladder”-shaped cluster states and focus on the non-classical correlations that develop between two qubits in the last rung of the ladder as measurements are done on other qubits in the cluster.
@@
In [25] we investigated the relationship between non-classical correlations in a subsystem of a globally entangled state and genuine multipartite entanglement in the global state. We found that when the global state of a collection of qubits considered has a regular structure like a cluster state, the two are typically related. Curiously, precisely when the the global state is a stabilizer state, the subsystem discord vanishes even if the global state is highly entangled. We take this line of investigation forward here, but in a slightly different direction. The ladder state is initially a stabilizer and the two qubits in the last rung share no quantum discord among them. We look at the non-classical correlations that develop between these two qubits as all the qubits in the cluster except the last three are measured. The measurements on the remaining qubits are chosen such that after each measurement, the state of the remaining qubits is pushed further away from being a stabilizer state. We choose to focus on the last three qubits of the ladder and study the quantum discord and entanglement between two of them because if the third one is also measured, the state of the remaining two will be pure and entanglement will be the only form of non-classical correlations present. With the third qubit also left unmeasured, there can be other forms of non-classical correlations between the last two.
@@
We can also view the last two qubits as an open quantum system coupled to a rather idealised environment that forms the rest of the ladder-like cluster state. From this perspective, we see how non-classical correlations can be built up in the open system due to measurements of a random nature happening on the environment that it is coupled to. As suggested in [25], the presence of non-classical correlations in a mixed state quantum computer can then be viewed as a consequence of computationally useful (non-stabilizer) entanglement present in the pure state of a larger system into which the mixed state of the computer is embedded in. The simple ladder-state model allows us to explore quantitatively how such correlations change in response to random measurements occurring on the environment.
@@
This paper is organized as follows. In Sec. II we review cluster states and their properties including non-classical correlations in subsystems. Flow of non-classical correlations in the ladder state model is discussed in the next section. Our conclusions are in Sec. IV.
@@
A cluster state [26] can be defined on a simple, un-directed graphs in any dimension. Qubits in the state $|+\rangle=(|0\rangle+|1\rangle)\sqrt{2}$ are placed on each of the vertices and each of the edges correspond to a two-qubit entangling gate $U_{kl}$ performed on the qubits placed on the vertices, $k$ and $l$ connected by the edge. The entangling gate used is typically the Controlled-Z (CZ) gate. Cluster states are a special case of the more generic graph states [27, 28] wherein the initial states of the qubits on the vertices as well as the entangling operations done across the edges need not be uniform.
A cluster state can be constructed from a collection of qubits all in the state $|0\rangle$ through the application of the Hadamard and CZ gates only. Since these operations form a subset of the Clifford group [29] of operations on a collection of qubits, the resultant cluster state is a stabilizer state [21]. A stabilizer state $|\psi_{\rm stab}\rangle$ on $n$ qubits is the simultaneous eigenstate with $+1$ eigenvalue of $2^{n}$ operators, all of which are strings of Pauli operators (elements of the Clifford group). This allows for a succinct description of the state in terms of $n$ generators of the $2^{n}$ operators. Application of operators belonging to the Clifford group on such states results in states whose stabiliser generators can be computed directly from the generators of the initial state. In other words, Clifford group operations on stabilizer states produce other stabilizer states and any quantum information processing task restricted to stabilizer states and Clifford group operations can be efficiently simulated using classical means, as proven in the celebrated Gottesman-Knill theorem [30, 24].
@@ II.1 Measurement based quantum computing
MBQC represents a distinctive approach to quantum computation that diverges from the traditional gate-based model. Gate-based quantum computing proceeds via application of quantum gates acting on a register of qubits, thereby evolving the state of the register towards a designated final state from which the output of the computation can be obtained [31, 32]. MBQC, however, takes a different path that rely on measurements for implementing the quantum computational paradigm. Instead of applying quantum gates directly, MBQC utilizes a cluster state as the substrate for its operations [33]. Cluster states are an ideal starting point in that it provides a high degree of quantum entanglement to be used up as a resource as the computation proceeds. Given the graph corresponding to the desired cluster, the steps to be followed to prepare the cluster state is straightforward as well even if practical implementation can be challenging [34, 35]. Entanglement in a cluster state is remarkably resilient against local operations, making it more challenging to disrupt than other highly entangled states like Greenberger-Horne-Zeilinger (GHZ) states. In fact, a single local operation cannot annihilate all the entanglement in a multipartite cluster state, while the same is not true for a GHZ state [26, 36].
MBQC is typically implemented on cluster states defined by simple and regular graphs like a two dimensional square lattice. Clusters defined on 2D triagonal, hexagonal and Kagome lattices are also known to be universal resource states for MBQC [37]. In fact, highly entangled states that do not necessarily share the symmetry and robustness of entanglement of the clusters defined on these simple lattices are known to be not useful for MBQC [38, 39]. Starting from simple graphs like a 2D square lattice has the added advantage that given any quantum computation implemented through a gate-based quantum circuit, the same circuit can literally be “drawn” on a suitably large cluster state showing that both models are equivalent. Drawing such a circuit involves removal of several qubits from the cluster which is achieved by measuring all of them on to the eigenbasis of the Pauli-$Z$ operator (computational basis). Using simple and symmetric graphs for building the cluster is also helpful for writing down the stabilizer generators for the cluster. The $n$ generators are identified by associating a string of Pauli operators to each of the $n$ qubits in the lattice which consists of the Pauli-$X$ operator on the qubit along with $Z$ operators on all the qubits connected to it.
@@
Two measurement outcomes $s_{j}\in\{0,1\}$ are possible for each qubit and typically the measurement on the next qubit depends on the result of the measurement outcome on the previous one. Single qubit gates are implemented by measuring a row of connected qubits with suitable measurement angles $\varphi_{j}$. Up to four such measurements may be needed to implement an arbitrary single qubit rotation. A two qubit gate like the cnot gate can be implemented on the two-dimensional cluster through a similar sequence of measurements on two rows of qubits [21]. In short, through the choice of suitable measurements a complete set of universal quantum gates can be realized within MBQC. Measurements on an $n$-qubit cluster state reproduces the action of a gate-based quantum computation on a register of $d$ qubits in the sense that after $n-d$ measurements, the state of the remaining qubits of the cluster will match the output of the gate-based circuit up to irrelevant overall phases. Measurement of these $d$ output qubits in a standard basis produces the measurement statistics which form the (classical) output of the computation.
@@ III Non-Classical Correlations in MBQC using Cluster States
The output qubits have to be driven to a specific state with its own distinctive quantum and classical correlations by the sequence of measurements in the MBQC model. Our interest is in how such correlations develop between the output qubits. For investigating this problem we consider a simplified cluster state defined on a square lattice with a ladder shape as shown in Fig. 1. The two qubits on the right end of the ladder are the output qubits. When the ladder shaped cluster is in the initial stabilizer state, non-classical correlations are not present in the state of the two output qubits [25]. The question of interest is how such correlations develop as measurements are done on other qubits in the cluster leading to the localisation of the relevant information on these two qubits as the output of the computation. Do such correlations develop only when the immediate neighbours are measured or do they develop earlier indicating a flow of the quantum information delocalized across the cluster into these qubits at each stage? What is the nature of the non-classical correlations developed on the output qubits and how does it depend on the details of the measurements performed on the rest of the cluster. These are some of the questions that we investigate in the following.
@@
We quantify the non-classical correlations in $\rho_{34}^{\rm out}$ using quantum discord [40, 41, 42, 43, 44]. Quantum discord is essentially the difference between the total correlations and the classical correlations in a bipartite quantum state $\rho_{\rm AB}$ of two systems $A$ and $B$. It is an entropic measure and the total correlations are quantified in terms of the mutual information $I(A:B)=S(A)+S(B)-S(AB)$ between $A$ and $B$ with $S(X)=-{\rm tr}(\rho_{X}\log\rho_{X})$ denoting the vonNeumann entropy. Classical correlations are quantified in terms of the conditional entropy $S(A|\rho_{B}^{j})$ where $\rho_{B}^{j}$ is the post-measurement state of $B$ obtained when a measurement indexed by $j$ is performed on subsystem $B$. In terms of this, an alternate expression for the mutual information can be written down as $J(A:B)=S(A)-\sum_{j}S(A|\rho_{B}^{j})$, where the sum over $j$ ensures a complete set of measurements on $B$. If $A$ and $B$ are classical systems with the vonNeumann entropy replaced by the corresponding Shannon entropy, then by Bayes theorem, the two definitions of classical mutual information $I$ and $J$ are equal to each other. In the quantum case the two need not be equal. Defining ${\mathcal{J}}(A:\bar{B})=\max_{\{\Pi_{j}\}}J(A|B)$, where the maximum is over all possible measurements $\{\Pi_{j}\}$ on $B$, quantum discord is defined as the difference,
@@
We see that the two generators $g_{4}(2)$ and $g_{4}(4)$ remain Pauli strings, but the generator $g_{4}(3)$ is not a Pauli string anymore but is a sum of two Pauli strings instead. In other words, the state of qubits $2$, $3$ and $4$ is no longer a stabilizer state. Since qubit $4$ is not directly connected to the measured qubit, the generator $g_{4}(4)$ associated with it remains invariant. At $\varphi_{1}=0,\pi/2,3\pi/2,$ and $2\pi$, the generator $g_{4}(3)$ also becomes a single Pauli string and for these values of $\varphi_{1}$, the quantum discord $D_{3\bar{4}}$ is absent.
@@ III.1 Six and eight qubit ladders
Figure 4: The top panel shows $D_{\bar{5}6}$ as a function of $\varphi_{1}$ for 20,000 random choices of $(\varphi_{1},\varphi_{2},\varphi_{3})$ for the six qubit ladder state. The middle and bottom panel shows $D_{\bar{5}6}$ as function of $\varphi_{2}$ and $\varphi_{3}$ respectively. $D_{\bar{5}6}$ for each choice of $(\varphi_{1},\varphi_{2},\varphi_{3})$ are averaged over the eight possible measurement outcomes $(s_{j}=\pm 1)$.
We can track the flow of non-classical correlations in the six qubit ladder by computing $D_{\bar{5}6}$ when measurements are done one by one on qubits $1$, $2$ and $3$. When the measurements are on the first and second qubits only then both $D_{5\bar{6}}$ and $D_{\bar{5}6}$ are zero. This is because the measurement operators commute with the CZ gates applied between qubits (3,5), (4,6), and (5,6), thereby leaving the output state (5,6) unaffected. However, the effect of the measurements on $1$ and $2$ do eventually flow to qubits $5$ and $6$ when qubit $3$ is measured as can be seen from Fig. 4. As a function of the last measurement angle $\varphi_{3}$ shown in the bottom panel of Fig. 4, $D_{\bar{5}6}$ shows a discernible pattern with two distinct oscillatory patterns. As a function of the second-last measurement angle $\varphi_{2}$, the same pattern persists but is a bit less well defined. The top panel of Fig. 4 shows that even if there is a pattern in the relationship between $D_{\bar{5}6}$ and the first measurement angle $\varphi_{1}$, it is masked by the subsequent measurements and does not show up as clearly as in the other case. Note that the maximum values for $D_{\bar{5}6}$ for the six qubit case is same as the maximum for $D_{3\bar{4}}$ for four qubits. The stabilizer generators $g_{6}(4)$, $g_{6}(5)$ and $g_{6}(6)$ transform under the measurements specified by $(\varphi_{1},\varphi_{2},\varphi_{3})$ as
@@
Quantum discords $D_{7\bar{8}}$ and $D_{\bar{7}8}$ averaged over the $322^{5}=$ 32 possible measurement outcomes on the proceeding 5 qubits are plotted in Fig. 6 against the angle $\varphi_{5}$ of the measurement on qubit $5$ which is directly connected to qubit $7$ of the cluster. The discord is computed again for 20,000 randomly chosen values of the measurement angles $\varphi_{i}$, $i=1,\ldots,5$. We see that with the longer ladder, the non-classical correlations developed between the qubits in the last rung is stronger with both $D_{7\bar{8}}$ and $D_{\bar{7}8}$ having a maximum value of $1$. The asymmetry between the two discords also is not present indicating the presence of entanglement between the two. Entanglement as quantified by the concurrence was also computed for the 20,000 choices of measurement angles and the result is plotted in Fig. 7 showing that indeed the qubits in the last rung of the ladder does get entangled when the ladder is sufficiently long.